OpenStudy (mckenzieandjesus):

Graph the data in the table below. Which kind of function best models the data? Write an equation to model the data. quadratic; y = –x2 + 2x – 3 linear; y = – 2x – 3 exponential; y = –3•1.5x quadratic; y = –x2 – 3

OpenStudy (mckenzieandjesus):



OpenStudy (mckenzieandjesus):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far (in terms of work)?

OpenStudy (mckenzieandjesus):

I dont understand how to do it

jimthompson5910 (jim_thompson5910):

ok let's say we pick on choice B linear; y = – 2x – 3

jimthompson5910 (jim_thompson5910):

IF the answer was choice B, then (0,-3) should lie on the graph so should (1,-2) and (2,-3) etc etc

jimthompson5910 (jim_thompson5910):

but what is the value of y when x = 1 y = – 2x – 3 y = -2*1 - 3 y = ??

OpenStudy (mckenzieandjesus):

-5?

jimthompson5910 (jim_thompson5910):

yes so (1,-5) is actually on y = – 2x – 3 not (1,-2)

jimthompson5910 (jim_thompson5910):

this eliminates choice B

OpenStudy (mckenzieandjesus):

ok

jimthompson5910 (jim_thompson5910):

test the points with the other equations. You need to make sure all points fit

OpenStudy (mckenzieandjesus):

so would i do -2*2-3?

OpenStudy (mckenzieandjesus):

nvrm!

jimthompson5910 (jim_thompson5910):

well you could, but we proved B wasn't the answer. So there's no need to check B any more

OpenStudy (mckenzieandjesus):

got it nvrm lol

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (mckenzieandjesus):

Would the answer be A?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (mckenzieandjesus):

Thank You!. I never understood those but now I do. A lot simpler than I thought.